Tamarisk, Inc. purchased a delivery truck for $29,200 on January 1, 2019. The truck has an expected salvage value of $2,200, and is expected to be driven 100,000 miles over its estimated useful life of 8 years. Actual miles driven were 16,100 in 2019 and 12,800 in 2020.

A. Calculate depreciable cost per mile under units-of-activity method.

B.Compute depreciation expense for 2019 and 2020 using (1) the straight-line method, (2) the units-of-activity method, and (3) the double-declining-balance method. (Round depreciable cost per unit to 2 decimal places, e.g. 0.50 and depreciation rate to 0 decimal places, e.g. 15%. Round final answers to 0 decimal places, e.g. 2,125.)

Depreciation Expense

2019

2020

(1) Straight-line method $enter a dollar amount rounded to 0 decimal places

$enter a dollar amount rounded to 0 decimal places

(2) Units-of-activity method $enter a dollar amount rounded to 0 decimal places

$enter a dollar amount rounded to 0 decimal places

(3) Double-declining-balance method $enter a dollar amount rounded to 0 decimal places

$enter a dollar amount rounded to 0 decimal places

C.

Assume that Tamarisk uses the straight-line method. Prepare the journal entry to record 2019 depreciation. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Round answers to 0 decimal places, e.g. 2,125.)

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount

D. Assume that Tamarisk uses the straight-line method. Show how the truck would be reported in the December 31, 2019, balance sheet. (Round answers to 0 decimal places, e.g. 2,125.)
Tamarisk, Inc.
Partial Balance Sheet
choose the accounting period

For the Month Ended December 31, 2019December 31, 2019For the Year Ended December 31, 2019

enter a balance sheet item

$enter a dollar amount

select between addition and deduction

AddLess

: enter a balance sheet item

enter a dollar amount

$enter a total of the two previous amounts
Answer #1

A. Calculate depreciable cost per mile under units-of-activity method.

Depreciation cost per mile = (29200-2200)/100000 = 0.27 per mile

B.Compute depreciation expense for 2019 and 2020 using (1) the straight-line method, (2) the units-of-activity method, and (3) the double-declining-balance method. (Round depreciable cost per unit to 2 decimal places, e.g. 0.50 and depreciation rate to 0 decimal places, e.g. 15%. Round final answers to 0 decimal places, e.g. 2,125.)

Depreciation Expense

2019

2020

(1) Straight-line method (29200-2200)/8 = 3375 3375
(2) Units-of-activity method 16100*.27 = 4347 12800*.27 = 3456
(3) Double-declining-balance method 29200*25% = 7300 29200*75%*25% = 5475

C.

Assume that Tamarisk uses the straight-line method. Prepare the journal entry to record 2019 depreciation. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Round answers to 0 decimal places, e.g. 2,125.)

Account Titles and Explanation

Debit

Credit

Depreciation expense 3375
Accumulated depreciation-delivery truck 3375
D. Assume that Tamarisk uses the straight-line method. Show how the truck would be reported in the December 31, 2019, balance sheet. (Round answers to 0 decimal places, e.g. 2,125.)
Tamarisk, Inc.
Partial Balance Sheet
December 31, 2019
Delivery truck 29200
Less: Accumulated depreciation -3375
$25825